### v1

@@ 3.3 Disparity in Transaction Execution
Having established that blocks by different builders are different, we now turn to study whether the execution of a transaction changes depending on which builder includes it in a block. For instance, a token swap on a DEX may yield different outcomes depending on the transactions that precede it. We focus on swaps on major DEX protocols — Uniswap V2, Uniswap V3, Pancakeswap, and Sushiswap — that are included in multiple blocks. We compare their execution across blocks using transaction logs, which capture the sequence of contract calls and resulting outputs. This enables us to detect differences in how the same transaction is executed depending on block context.
@@
The first dimension along which a transaction’s execution may vary depending on the block in which it is included is its execution speed. We identify 12 user swaps that first appeared during the bidding cycle for block 21,322,649 and were still present in the subsequent cycle. Of these 12 delayed transactions, 5 were exclusive to Titan during the 21,322,649 cycle. The remaining 7 were present in blocks by Flashbots, Buildernet, and Titan. [^15] These 7 transactions are different from the 3 transactions identified in the previous subsection because these 3 transactions were present only during the bidding cycle for block 21,322,649. In the following cycle, 3 of the 5 transactions initially exclusive to Titan remain exclusive, while the other 2 change status to “private but not exclusive,” having been submitted via both MEV Blocker and Flashbots Protect. All 9 transactions that were not exclusive to Titan in the second cycle are included in the winning block for that cycle (21,322,650), which was built by Beaverbuild. 2 of the remaining 3 transactions are included in block 21,322,651 (built by Titan), for which we do not have complete bidding data. [^16] The hashes of those transactions are
@@
The last transaction is a user’s swap that remains exclusive to Titan and included in winning block 21,322,652 (also built by Titan), but fails. [^17] Its hash is 0x74b6b90aa9ec32adc9e35c15fbd1cf5af37a8ee4d43141785ed21f3bef3f6d8b These findings suggest that which builder has access to which transactions affects not only whether transactions are included on-chain, but also when they are included.
The second dimension along which a transaction’s execution may vary depending on the block in which it is included is whether the transaction succeeds or fails. For the auction cycle under consideration, we identify seven transactions that fail in at least one block: six of these fail in all blocks in which they appear, and five are included in the winning block. Among the six transactions that always fail, four also fail when we simulate them at the top of the block in which they were included, suggesting that they were erroneously constructed or misconfigured. At least for this bidding cycle, whether a swap fails does not seem to depend on the block in which it was included.
Finally, the same swap transaction may be executed at different prices depending on the block in which it is included. Among the 12 users’ swaps observed across multiple bidding cycles, only 3 exhibit variation in execution price. Surprisingly, for those 3 swaps, execution quality is better in the later cycles. We also identify 7 users’ transactions with different execution prices across blocks within the bidding cycle for block 21,322,649. Two swap ETH for MogCoin and receive the best execution in a block built by Rsync, where they are preceded by a swap by Rsync-bot selling MogCoin. However, we also find 3 transactions selling MogCoin that would have achieved better execution in blocks built by non-searcher-integrated builders (i.e., neither Rsync nor Titan). In the remaining 2 cases, better execution is also offered by non-integrated builders. Overall, transactions that move in the opposite direction of searchers’ tend to receive better execution in blocks built by searcher-integrated builders, while those swapping in the same direction of searchers’ perform better in blocks built by non-integrated builders.
@@
Having established that the identity of the winning builder affects the quality of users’ transaction execution, we now turn to the behavior of Rsync-bot and Titan-bot. As noted earlier, during this auction cycle, these two bots account for between 94% and 97% of the total value of submitted blocks, making them one of the main drivers of the auction’s outcome.
@@
    the fees each bot pays for block inclusion and how those fees change during the auction cycle;
@@
Furthermore, because both searchers are integrated with their respective builders, we can interpret their willingness to pay for block inclusion as the expected profits from the arbitrage trade. This, in turn, allows us to infer a “risk-adjusted implied price” on the centralized exchange (CEX) side of the arbitrage path—a component that is typically unobservable. This method provides new insights into the off-chain leg of arbitrage activity.
@@
Note that all prices here are quoted in ETH. In some cases, such as trades involving stablecoins (e.g., USDC or USDT), we may prefer to express prices in terms of the other token. This simply requires inverting the expressions above.
@@
Finally, we adjust the base currency in the graphs for the implied centralized exchange (CEX) price so that higher prices are always better for the bots. The main result here is that, for the MOG/WETH arbitrage, Titan-bot’s implied CEX price is 1.9% better than that of Rsync-bot (Note that MOG Coin was not listed on Binance nor Coibase during our study period). For USDT/WETH and USDC/WETH, instead, Rsync-bot seems to have an advantage of approximately 0.17% and 0.02%, respectively. Note also that the implied CEX prices are sometimes above or below Binance’s. Remember that the implied CEX prices are risk-adjusted, that is, the arbitrage risk is reflected in a less favorable implied CEX price. Also, the reported Binance price does not account for trading fees or price impact. These reasons may explain why the bots’ implied CEX prices may be below those on Binance. At the same time, the bots may have sources of liquidity other than Binance or some inventory, which they may use to achieve an even better execution price than Binance’s.
@@ 4 Entire Dataset
(a) 11 seconds
(b) 12 seconds
(c) 13 seconds
(d) 14 seconds
Figure 8: Proportion of transactions shared between blocks by different builders, at different moments in the bidding cycle.
We now extend our analysis to all bidding cycles in the dataset. Figure 8 confirms the earlier pattern: blocks proposed by different builders differ in their transaction content, and this difference becomes more pronounced as the bidding cycle progresses.
We also investigate whether transactions are shared among builders. Again, we limit our analysis to transactions shared with Rsync or Titan. For each new transaction added by Rsync or Titan, we check whether it appeared in a block submitted by another builder at least one second earlier. [^18] Again, a builder’s new transactions are those included in a block submitted at least one second after a builder’s initial submission, and not present in that first block. We identify 20 such cases: 3 involving Rsync and 17 involving Titan. Of the 3 Rsync cases, two transactions were initially exclusive to Beaverbuild and were later included in a block by Rsync, which ultimately won. The third, a searcher’s swap, appeared in blocks from five different builders, but only entered a Rsync block four seconds after its first appearance (again, Rsync won the block). Of the 17 transactions later included by Titan, 8 were submitted via the Flashbots private mempool. Five originated from a single searcher address, 0x89a99a0a17d37419f99cf8dc5ffa578f3cdb58b5, and appear in Flashbots Protect’s dataset as shared with more than 15 builders. These transactions remained exclusive to Flashbots Builder for approximately 2.6 seconds before being included in Titan’s blocks. Another five transactions from the same address follow the same pattern (initial exclusivity to Flashbots followed by appearance in Titan’s blocks) but were not included on-chain. It is plausible that they were also shared via Flashbots Protect but were excluded from the dataset, which only logs transactions that are eventually included.
In summary, transaction sharing between builders during the bidding cycle does occur, but is relatively rare. In most cases, it involves transactions shared via Flashbots Protect and exclusive to Flashbots builder during the initial part of the bidding cycle.
@@ 4.1 Time of inclusion in a winning block
We identified 1,288 transactions that appeared in multiple bidding cycles. These transactions appeared in a bidding cycle before the winning block was selected, were not included in the winning block, and appeared again in the subsequent bidding cycle. Among these, 871 were present for two cycles, 198 for three cycles, and 108 for four cycles before being included in a winning block. Remarkably, one transaction remained present for 2,437 bidding cycles before inclusion; it was exclusive to the builder payload.de and was ultimately included only when payload.de won. In addition, 28 transactions appeared in multiple cycles but were never included in a winning block.
Among the transactions present in multiple bidding cycles, only 17 are searchers’ transactions. Of the remaining transactions, 701 first appeared in our dataset as public, 189 as private, and 381 as exclusive (with 374 exclusive to Titan Builder). inlineinlinetodo: inlineMaybe for another paper, but why are so many public transactions not picked up? Is the public mempool being ignored? Have private mempools crowded out public mempools? Has EIP -1559 failed, meaning: it is not enough to pay base fee + something to be included, but, again, users need to strategize on how much to pay? Transactions that are initially public remain public throughout all subsequent bidding cycles, and very few (13) transactions that begin as private later appear in the public mempool. However, the majority of exclusive transactions present over multiple bidding cycles change status over time. Specifically, 162 exclusive transactions remain exclusive across all cycles in which they appear, while 174 exclusive transactions transition to private.  [^19] We find a further 33 transactions that were initially exclusive to a builder and were included in a later auction cycle by that same builder. However, the auction cycle in which these transactions were included is not in our dataset. Hence, we cannot determine whether these transactions remained exclusive to the builder or transitioned to private. We also identify 12 transactions that were initially exclusive but later appeared in the public mempool. Of these, 9 were present during two bidding cycles, and 3 were present during three cycles. All are user transactions: three interacted with a DEX via the 0x router, and four via the 1inch router. [^20] These transactions are: 0x1daea1584d684385fb25209bad3a49c54a1e2c500e27543334c7b3e695a12ffb, 0x6605c1c36fcd3d696844519d6ff7199fe4ddd34fb4f3e0ed4da71aedfcf1c633, 0x3dadbab8e85b6e48a2c33f9029ea4ebebdd3def8b9d907faff15ef81527de968, 0x9c7364eddb94fe659bc8104585ff86a8a234480fe00a9f9aa3eb410e572ebdf2, 0x9d476dbc524651cbe26bcafab99f6fe40a823a5588fa5cd5002e5b414a8a2f8c, 0x417440b083bdcb529663c6615c32d89a92beb7863f1a5d6b12bca5b94f4a0f87, 0xe56caee27e79dee8a4c7ed6f945ccbc2cc733d761e19bd68329dc364c61e73a7, 0xae5b88bc0bffa04811eda0dfdcd7224599e2757339e18b2d57f4f40d61902a4c, 0xd41139b67a193045b4d60266c084f6dadd0cb95dccd7b2a06ea68653d6a313ec, 0x597fc0cb0610d1e3ea2baf0f6f90d0d741a6429bfdd9b375bfd2493b70d41dcd, 0x44e8105d5c672195877dc1ed970511e0f0a15fc25cbcafe71d51d428d54c215d, 0xd74718e6d8d58bc33334c0c2f6642d9dec49496329329d8fdcb1e698e81abc94.
To summarize: the vast majority of searchers’ transactions appear in only a single bidding cycle and are either included in the winning block or discarded. Hence, transactions present across multiple bidding cycles are predominantly users’ transactions. Given that our dataset includes 5,947 user transactions, we estimate that approximately 21% users’ transactions are delayed. This delay is partly explained by the fact that 30% of these transactions are, at least initially, exclusive to a single builder. This is surprising, as users should submit transactions privately to multiple builders or even broadcast them publicly (in case of simple transfers). Even more startling is that some initially exclusive transactions appear in the public mempool after one or two bidding cycles. These transactions are first delayed and then exposed to various attacks, which is clearly not in the users’ best interest.
@@ 4.2 Successful execution
A transaction may succeed or fail depending on the block in which it is included. To explore this possibility, we consider each transaction hash/block combination and classify it as a fail if its simulation outcome is status = 0, or if it produces no transaction log despite the same transaction generating one in another block. A success is defined as any transaction/block pair not classified as a failure. Under this definition, “failures” include both outright failures and reverts, such as a swap reverting because its execution price would be below the user’s specified worst acceptable price. In contrast, a “success” indicates that the transaction would have executed as intended, had the corresponding block been added to the chain. Among the 49 transactions that sometimes fail and sometimes succeed, 44 are swaps. These account for 318,742 transaction-block observations. We therefore focus our analysis exclusively on swaps.
@@
We also include transaction fixed effects: that is, a dummy variable for each unique transaction hash. These control for transaction-specific characteristics and capture the average likelihood that a transaction executes successfully across all blocks in which it appears. The coefficients on the other variables then measure how execution probability varies depending on block timing, builder identity, transaction ordering, and the presence of searcher bots.
@@
Table 1 reports three regressions, each on a different samples. The first regression yields mixed results: being included in a block built by Titan or Rsync reduces the probability of successful execution, but the presence of transactions from the corresponding bot increases it. In contrast, the second and third regressions reveal a clearer pattern. Swaps in the same direction as the bots are significantly more likely to fail when they are included in blocks built by Rsync or Titan, especially if those blocks also contain transactions from the bots themselves. The opposite holds for swaps in the opposite direction: their likelihood of success increases under the same conditions, although in this case the main effect is due to the presence of the bots. We also observe that swaps included later in a block (i.e., with a higher transaction index) are slightly less likely to fail. This result is, however, difficult to interpret, as transaction ordering within a block is determined by the builder. Finally, transactions are more likely to fail when included in blocks submitted later in the bidding cycle.
Using the estimated coefficients, we perform back-of-the-envelope calculations to quantify the effects. Compared to inclusion in a block built by a non-searcher-integrated builder, a swap in the same direction as Titan-bot or Rsync-bot is approximately 18% less likely to execute successfully when included in a block built by Titan or Rsync that also contains a transaction from the corresponding bot. By contrast, a swap in the opposite direction is approximately 1% more likely to execute successfully under the same conditions.
@@ 4.3 Execution price
The way a swap is included in a block may also influence its execution price. To investigate this possibility, we examine all swap in blocks where they executed successfully and the execution price can be computed. This yields 272 unique transaction hashes, 112,747 transaction-block observations, and 342,528 swap-block observations, which will be our unit of analysis (note that many transactions perform multiple swaps). inlineinlinetodo: inlineIt is kind of interesting that the majority of transactions performing swaps perform more than 1 swap
@@
where $p$ is the execution price of the swap in a given block. We adjust the base currency so that higher values of $p_{\text{norm}}$ always reflect better execution from the user’s perspective. The distribution of $p_{\text{norm}}$ has a median of $-1.41\times 10^{-14}$, with a minimum of $-10.53$ and a maximum of $13.47$ basis points.
@@
inlineinlinetodo: inlineAt some point we should try to increase the R2 of the regression - that is, we should find out what other variable has explanatory power on $p_{norm}$
@@ 4.4 Competition between Searchers
In the previous section, we highlighted the importance of the competition between Rsync-bot and Titan-bot. It turns out that these two searchers compete intensely not only during the auction cycle of winning block 21322649, but throughout our entire dataset. inlineinlinetodo: inlineThere is a bit of a gap in our reasoning here: in the previous section, we showed that those two bots are responsible for the vast majority of the value of the blocks, which is why we study them. Can we compute the same for the entire dataset? Are these two bots ”important” also when looking at all bidding cycles?
@@
We identify 57 instances in which Titan-bot and Rsync-bot compete for execution in the same DEX pool during the same auction cycle. In 52 cases, one of the tokens swapped is WETH, allowing us to apply the same methodology introduced in the previous section — namely, deriving an implied risk-adjusted CEX price for each bot. To enable direct comparison with observed market data, we further restrict the sample to tokens traded on Binance. This yields 29 auction cycle/DEX pool combinations for which we can compare the bots’ implied prices to contemporaneous Binance prices. We plot the implied DEX price alongside the Binance price for all 29 cases in Appendix 0.A. Note that Binance reports prices at one-second intervals. As expected, the implied DEX prices generally track the Binance price closely. This correlation is particularly strong in cases involving the USDT/WETH and USDC/WETH pairs, which account for the majority of the sample. A weaker but still visible relationship is observed for the UNI/WETH and AAVE/WETH pairs. In contrast, for the DAI/WETH pair, the implied DEX and Binance prices do not exhibit a consistent relationship, likely because Binance is not the primary centralized exchange for DAI trading. This is confirmed by the fact that the total volume on the DAI/ETH market on Binance during our study period was only 2.5 ETH. [^22] For comparison, total volume on Binance during the study period for the other markets was: 1,487 ETH for the USDC/ETH market, 16,209 ETH for the USDT/ETH market, 1,374 ETH for the WBTC/ETH market, 6.9 ETH for the UNI/ETH market, 36 ETH for the AAVE/ETH Binance market.
@@
We complement the visual analysis with a regression approach. For each second $s$, we compute the average implied CEX price and trade volume for Rsync-bot and Titan-bot by aggregating all trades executed by each bot between seconds $s-1$ and $s$. We denote the average implied CEX price by $p_{\text{implied}}$, and associate each observation with timestamp $s$11todo: 1It may be more accurate to center the timestamp at $s-0.5$ or $s+0.5$; we leave this refinement for future work. For technical reasons, we restrict our analysis to the USDT/WETH and USDC/WETH trading pairs. [^23] As discussed earlier, Binance prices for some token pairs appear unreliable. More importantly, the number of observations, values of $p_{\text{implied}}$, and trade volumes vary considerably across tokens. We have 104 observations for WETH/USDC and WETH/USDT combined, but only 24 for WETH/DAI, 18 for WETH/UNI, and 5 or fewer for WBTC/WETH and AAVE/WETH. Moreover, the average value of $p_{\text{implied}}$ is around 3,500 for stablecoin pairs (USDC, USDT, DAI), but drops to 261 for UNI/WETH, 15 for AAVE/WETH, and 0.037 for WBTC/WETH. Trade volumes also vary: average volume is 320.3 ETH for WETH/WBTC, compared to 43.1 ETH for other assets. As a result, separate regressions on the less common pairs are underpowered, while pooled regressions suffer from multicollinearity due to token-specific heterogeneity. This filtering results in 104 total observations: 53 for Rsync-bot and 51 for Titan-bot. We regress $p_{\text{implied}}$ on the contemporaneous Binance price, as well as its one-second lag and lead. The regression results, shown in Table 3, indicate that only the contemporaneous and one-second lagged Binance prices are significantly correlated with the bots’ implied CEX prices. Among these, the contemporaneous price has the strongest explanatory power. Therefore, the contemporaneous Binance price serves as the most appropriate benchmark for evaluating $p_{\text{implied}}$.
@@
Binance Price (+0 sec.) $0.6587$*** $0.106$
Binance Price (+1 sec.) $-0.1197$ $0.128$
Binance Price (+2 sec.) $-0.1306$ $0.095$
Binance Price (-1 sec.) $0.4783$*** $0.097$
Binance Price (-2 sec.) $0.1001$ $0.060$
Constant $46.192$* $19.030$
Observations 104
R-squared 0.997
@@
In both cases, a higher $p_{\text{diff}}$ indicates better execution than Binance. We then calculate
@@
as the percentage price improvement relative to the contemporaneous Binance price. The minimum, maximum, and average values of $p_{\text{Improvement}}$ are -0.017652, 0.064844, and 0.011934, respectively.
Table 4 reports the coefficients of two separate regressions. Both include a constant, a dummy variable equal to 1 if the observation is from Titan bot, and time measured in seconds since the last block. The difference between the two regressions is that the second includes swap volume (in WETH) in piecewise linear form, where each piece corresponds to 50 ETH. The coefficient on the constant represents the average percentage price improvement ($p_{\text{Improvement}}$) for a zero-volume swap by Rsync-bot, assuming the block was built at time zero in the bidding cycle. The coefficient on Is Titan bot captures the difference in execution quality between Titan-bot and Rsync-bot. The results indicate that the bots achieve better execution than the Binance price, and that Rsync-bot outperforms Titan-bot. Execution quality seems to decline with trade size, although this result needs to be taken with a grain of salt because it is driven by the only 2 trades of volume larger than 150. Using the estimated coefficients, we perform back-of-the-envelope calculations. For a swap of volume lower than 150 ETH the implied CEX price for ETH is approximately 4.2 basis points better than the contemporaneous Binance price for Rsync-bot, and 3.4 basis points better for Titan-bot.
@@
Time since Last Block -0.0015 -0.0019
@@
Volume X (Volume lower than 50) -0.0015
@@
Volume X (Volume between 50 and 100) -0.0043
@@
Volume X (Volume between 100 and 150) -0.0061
@@
Volume X (Volume greater than 150) -0.0417***
0.009
Constant 0.0368** 0.0420***

### v2

@@ 3.2 Disparity in Transaction Execution
We now turn to study whether the execution of a transaction changes depending on which builder includes it in a block. For instance, a token swap on a DEX may yield different outcomes depending on the transactions that precede it. We focus on swaps on major DEX protocols — Uniswap V2, Uniswap V3, Pancakeswap, and Sushiswap — that are included in multiple proposed blocks. We compare their execution across blocks using transaction logs, which capture the sequence of contract calls and resulting outputs. This enables us to detect differences in how the same transaction is executed in different blocks.
@@
The first dimension along which a transaction’s execution may vary depending on the block in which it is included is its execution speed. We identify 12 user swaps that first appeared during the bidding cycle for block 21,322,649 and were still present in the subsequent cycle. Of these 12 delayed transactions, 5 were exclusive to Titan during the 21,322,649 cycle, meaning that they appeared only in blocks built by Titan. The remaining 7 were present in blocks by Flashbots, Buildernet, and Titan. [^18] These 7 transactions are different from the 3 transactions identified in the previous subsection because these 3 transactions were present only during the bidding cycle for block 21,322,649. In the following cycle, 3 of the 5 transactions initially exclusive to Titan remain exclusive, while the other 2 appear in blocks by multiple builders, having been submitted via both MEV Blocker and Flashbots Protect. All 9 transactions that were not exclusive to Titan in the second cycle are included in the winning block for that cycle (21,322,650), which was built by Beaverbuild. 2 of the remaining 3 transactions are included in block 21,322,651 (built by Titan), for which we do not have data other than the winning block. [^19] The hashes of those transactions are
@@
The last transaction is a user’s swap that remains exclusive to Titan and included in winning block 21,322,652 (also built by Titan). [^20] Its hash is 0x74b6b90aa9ec32adc9e35c15fbd1cf5af37a8ee4d43141785ed21f3bef3f6d8b Hence which builder has access to which transactions affects not only whether transactions are included on-chain, but also when they are included.
The second dimension along which a transaction’s execution may vary depending on the block in which it is included is whether the transaction succeeds or fails. For the auction cycle under consideration, we identify seven transactions that fail in at least one block: six of these fail in all blocks in which they appear and four of those always fail also when we simulate them at the top of each block in which they were included, suggesting that they were erroneously constructed or misconfigured. At least for this bidding cycle, whether a swap fails does not seem to depend on the block in which it was included.
Finally, the same swap transaction may be executed at different prices depending on the block in which it is included. Among the 12 users’ swaps observed across multiple bidding cycles, only 3 exhibit variation in execution price. Surprisingly, for those 3 swaps, execution quality is better in the later cycles. We also identify 7 users’ transactions with different execution prices across blocks within the bidding cycle for block 21,322,649. Two swap ETH for MogCoin and receive the best execution in a block built by Rsync, where they are preceded by a swap by Rsync-bot selling MogCoin. However, we also find 3 transactions selling MogCoin that achieve worse execution in blocks built by Rsync and Titan (compared to other builders). Overall, transactions that trade in the opposite direction as the two bots tend to receive better execution in blocks built by Rsync and Titan, while those swapping in the same direction as the two bots receive worse execution in blocks built by Rsync and Titan.
@@
Having established that the identity of the winning builder affects the quality of users’ transaction execution, we now turn to the behavior of Rsync-bot and Titan-bot. As noted earlier, during this auction cycle, these two bots account for between 94% and 97% of the total revenues of submitted blocks, making them one of the main drivers of the auction’s outcome.
@@
    the fees each bot pays and how they change during the auction cycle;
@@
Furthermore, by interpreting the fee attached to each transaction as the expected profits in case that transaction is added on chain, we can infer a “risk-adjusted implied price” on the centralized exchange (CEX) side of the arbitrage path—a component that is typically unobservable. This method provides new insights into the off-chain leg of arbitrage activity.
@@
Note that all prices here are quoted in ETH. In some cases, such as trades involving stablecoins, we may prefer to express prices in terms of the other token. This simply requires inverting the expressions above.
@@
Finally, we adjust the base currency in the graphs for the implied centralized exchange (CEX) price so that higher prices are always better for the bots. The main result here is that, for the MOG/WETH arbitrage, Titan-bot’s implied CEX price is 1.9% better than that of Rsync-bot (Note that MOG Coin was not listed on Binance nor Coibase during our study period). For USDT/WETH and USDC/WETH, instead, Rsync-bot seems to have an advantage of approximately 0.17% and 0.02%, respectively. Note also that the implied CEX prices are sometimes above or below Binance’s. Remember that the implied CEX prices are risk-adjusted, that is, the arbitrage risk is reflected in a less favorable implied CEX price. Also, the reported Binance price does not account for trading fees or price impact. These reasons may explain why the bots’ implied CEX prices may be below those on Binance. At the same time, the bots may trade on multiple centralized markets or have their own inventory, which they may use to achieve an even better execution price than Binance’s.
@@ 4 Entire Dataset
We now extend our analysis to all bidding cycles in the dataset. We first examine differences in transaction execution across blocks and then analyze the behavior of Titan-bot and Rsync-bot.
@@ 4.1 Time of inclusion in a winning block
We identify 1,288 transactions that were present in multiple bidding cycles. These transactions appeared in a bidding cycle before the winning block was selected, were not included in the winning block, and appeared again in the subsequent bidding cycle. Among these, 871 were present for two cycles, 198 for three cycles, and 108 for four cycles before being included in a winning block. Remarkably, one transaction remained present for 2,437 bidding cycles before inclusion; it was exclusive to the builder payload.de and was ultimately included only when payload.de won. In addition, 28 transactions appeared in multiple cycles but were never included in a winning block. Among the transactions present in multiple bidding cycles, only 17 are searchers’ transactions. Of the remaining transactions, 701 first appeared in our dataset as public, 189 as private, and 381 as exclusive (with 374 exclusive to Titan Builder).
Hence, the vast majority of searchers’ transactions appear in only a single bidding cycle and are either included in the winning block or discarded. Transactions present across multiple bidding cycles are predominantly users’ transactions. Given that our dataset includes 5,947 user transactions, we estimate that approximately 21% users’ transactions are delayed. This delay is partly explained by the fact that 30% of these transactions are, at least initially, exclusive to a single builder who does not win. [^21] For a more in depth analysis, see the companion paper Pahari and Canidio (2025).
@@ 4.2 Successful execution
A transaction may succeed or fail depending on the block in which it is included. To explore this possibility, we consider each transaction hash/block combination and classify it as a fail if its simulation outcome is status = 0, or if it produces no transaction log despite the same transaction generating one in another block. A success is defined as any transaction/block pair not classified as a failure. Under this definition, “failures” include both outright failures and reverts, such as a swap reverting because its execution price would be below the user’s specified worst acceptable price. In contrast, a “success” indicates that the transaction would have executed as intended, had the corresponding block been added to the chain. Among the 49 transactions that sometimes fail and sometimes succeed, 44 are swaps. We therefore focus our analysis exclusively on swaps, which account for 318,742 transaction-block observations.
@@
We also include transaction fixed effects: that is, a dummy variable for each unique transaction hash. These control for transaction-specific characteristics and capture the average likelihood that a transaction executes successfully across all blocks in which it appears. The coefficients on the other variables then measure how the probability of successful execution varies depending on block timing, builder identity, transaction ordering, and the presence of the two bots.
@@
Table 1 reports three regressions, each on a different samples. The first regression yields mixed results: being included in a block built by Titan or Rsync reduces the probability of successful execution, but the presence of transactions from the corresponding bot increases it. In contrast, the second and third regressions reveal a clearer pattern. Swaps in the same direction as the bots are significantly more likely to fail when they are included in blocks built by Rsync or Titan, especially if those blocks also contain transactions from the bots themselves. The opposite holds for swaps in the opposite direction: their likelihood of success increases in blocks which also contain transactions from the bots. We also observe that swaps included later in a block (i.e., with a higher transaction index) are slightly less likely to fail. This result is, however, difficult to interpret, as transaction ordering within a block is determined by the builder. Finally, transactions are more likely to fail when included in blocks submitted later in the bidding cycle.
Using the estimated coefficients, we perform back-of-the-envelope calculations to quantify the effects. Compared to inclusion in a block built by a non-searcher-integrated builder, a swap in the same direction as Titan-bot or Rsync-bot is approximately 18% less likely to execute successfully in blocks that also contain a transaction from these bots (remember that almost all blocks with transactions from these bots are built by the corresponding builder). By contrast, a swap in the opposite direction is approximately 1% more likely to execute successfully under the same conditions.
@@ 4.3 Execution price
The way a swap is included in a block may also influence its execution price. To investigate this possibility, we examine all swap in blocks where they executed successfully and the execution price can be computed. This yields 272 unique transaction hashes, 112,747 transaction-block observations, and 342,528 swap-block observations, which will be our unit of analysis (note that many transactions perform multiple swaps).
@@
where $p$ is the execution price of the swap in a given block. We adjust the base currency so that higher values of $p_{\text{norm}}$ always reflect better execution from the user’s perspective. The distribution of $p_{\text{norm}}$ has a median of $1.41\times 10^{-14}$, a minimum of 10.53-10.53 and a maximum of $13.47$ basis points.
@@ 4.4 Competition between Searchers
In the previous section, we highlighted the importance of competition between Rsync-bot and Titan-bot. It turns out that these two searchers compete intensely throughout our entire study period. And, like in auction for block 21,322,649, their competition is a key determinant of which builder wins the block: every auction cycle for which we have data includes at least one transaction from these bots, which together account for approximately 60% of the builders’ revenues in the blocks where they appear, underscoring their central role in shaping auction outcomes.
@@
We identify 57 instances in which Titan-bot and Rsync-bot compete for execution in the same DEX pool during the same auction cycle. In 52 cases, one of the tokens swapped is WETH, allowing us to apply the same methodology introduced in the previous section — namely, deriving an implied risk-adjusted CEX price for each bot. To enable direct comparison with observed market data, we further restrict the sample to tokens traded on Binance. This yields 29 auction cycle/DEX pool combinations for which we can compare the bots’ implied prices to contemporaneous Binance prices. We plot the implied DEX price alongside the Binance price for all 29 cases in Appendix 0.A (note that Binance reports prices at one-second intervals, we use the opening price). As expected, the implied DEX prices generally track the Binance price closely. This correlation is particularly strong in cases involving the USDT/WETH and USDC/WETH pairs, which account for the majority of the sample. A weaker but still visible relationship is observed for the UNI/WETH and AAVE/WETH pairs. In contrast, for the DAI/WETH pair, the implied DEX and Binance prices do not exhibit a consistent relationship, likely because Binance is not the primary centralized exchange for DAI trading. This is confirmed by the fact that the total volume on the DAI/ETH market on Binance during our study period was only 2.5 ETH. [^23] For comparison, total volume on Binance during the study period for the other markets was: 1,487 ETH for the USDC/ETH market, 16,209 ETH for the USDT/ETH market, 1,374 ETH for the WBTC/ETH market, 6.9 ETH for the UNI/ETH market, 36 ETH for the AAVE/ETH Binance market.
@@
We complement the visual analysis with a regression approach. For each second $s$, we compute the average implied CEX price and trade volume for Rsync-bot and Titan-bot by considering all trades executed by each bot between seconds s0.5s-0.5 and $s+0.5$. We denote the average implied CEX price by $p_{\text{implied}}$, and associate each observation with timestamp $s$. For technical reasons, we restrict our analysis to the USDT/WETH and USDC/WETH trading pairs. [^24] As discussed earlier, Binance prices for some token pairs appear unreliable. More importantly, the number of observations, values of $p_{\text{implied}}$, and trade volumes vary considerably across tokens. We have 104 observations for WETH/USDC and WETH/USDT combined, but only 24 for WETH/DAI, 18 for WETH/UNI, and 5 or fewer for WBTC/WETH and AAVE/WETH. Moreover, the average value of $p_{\text{implied}}$ is around 3,500 for stablecoin pairs (USDC, USDT, DAI), but drops to 261 for UNI/WETH, 15 for AAVE/WETH, and 0.037 for WBTC/WETH. Trade volumes also vary: average volume is 320.3 ETH for WETH/WBTC, compared to 43.1 ETH for other assets. As a result, separate regressions on the less common pairs are underpowered, while pooled regressions suffer from multicollinearity due to token-specific heterogeneity. This filtering results in 89 total observations: 43 for Rsync-bot and 46 for Titan-bot. We regress $p_{\text{implied}}$ on the contemporaneous Binance opening price, as well as its one- and two-second lag and one-second lead. The regression results, shown in Table 3, indicate that only the contemporaneous and one-second lagged Binance prices are significantly correlated with the bots’ implied CEX prices. Among these, the contemporaneous price has the strongest explanatory power. Therefore, the contemporaneous Binance price serves as the most appropriate benchmark for evaluating $p_{\text{implied}}$.
@@
Binance Price (0 sec.) 0.79450.7945*** $0.108$
Binance Price (+1 sec.) 0.1491-0.1491 $0.129$
Binance Price (-1 sec.) 0.25130.2513** $0.108$
Binance Price (-2 sec.) $0.0972$ $0.063$
Constant $21.5424$ $23.070$
Observations 89
R-squared 0.998
@@
In both cases, a higher $p_{\text{diff}}$ indicates better execution for the bot. We then calculate
@@
as the percentage price improvement relative to the contemporaneous Binance price. The minimum, maximum, and average values of $p_{\text{Improvement}}$ are -0.01428, 0.04655 and 0.01285 respectively.
Table 4 reports the coefficients of two separate regressions. Both include a constant, a dummy variable equal to 1 if the observation is from Titan bot, and time (seconds) since the last block. The difference between the two regressions is that the second includes swap volume (in WETH) in piecewise linear form, where each piece corresponds to 50 ETH. The coefficient on the constant represents the average percentage price improvement ($p_{\text{Improvement}}$) for a zero-volume swap by Rsync-bot, assuming the block was built at time zero in the bidding cycle. The coefficient on Is Titan bot captures the difference in execution quality between Titan-bot and Rsync-bot. The results indicate that the bots achieve better execution than the Binance price, and that Rsync-bot outperforms Titan-bot. Execution quality seems to decline with trade size, although this result needs to be taken with a grain of salt because there are only 2 trades of volume larger than 150. Using the estimated coefficients, we perform back-of-the-envelope calculations. For a swap of volume lower than 100 ETH the implied CEX price for ETH is approximately 2.8 basis points better than the contemporaneous Binance price for both bots.
@@
Time since Last Block -0.009 -0.001
@@
Volume $\times$ (Volume lower than 50) -0.0007
@@
Volume $\times$ (Volume between 50 and 100) 0.0003
@@
Volume $\times$ (Volume between 100 and 150) 0.0066
@@
Volume $\times$ (Volume greater than 150) 0.0231***
0.008
Constant 0.028 0.0287**